OpenStudy (anonymous):

The sum of the squares of 3 consecutive positive integers is 116. What are the numbers? Which of the following equations is used in the process of solving this problem?

OpenStudy (anonymous):

Which of the following equations is used in the process of solving this problem? 3n2+ 5 = 116 3n2+ 3n + 3 = 116 3n2+ 6n + 5 = 116

zepdrix (zepdrix):

You understand what consecutive means, yes? :) So let's start by letting an unknown number, `the first of the 3 consecutive numbers`, be called \(\large n\). That means the next number will be \(\large n+1\). And the number after that \(\large n+2\). So these are our 3 consecutive numbers. ---------------------------------------------------------- The sum of their SQUARES is 116. So let's go ahead and write this out.\[\large n^2+(n+1)^2+(n+2)^2=116\]This is how we write out the sum of their squares. It's the first number squared + the second number squared + the third number squared.

zepdrix (zepdrix):

To match one of the forms listed, we'll expand the squares.\[\large n^2+(n^2+n+n+1)+(n^2+2n+2n+4)=116\]If you're confused about how I expanded those out, let me know. From here, just simplify! :)